# TimeChat-Captioner: Scripting Multi-Scene Videos with Time-Aware and Structural Audio-Visual Captions

### v1

@@ Abstract
Figure 1: Illustration of the OmniDenseCaptioning task. This paper introduces Omni Dense Captioning task, which generates fine-grained, temporally grounded descriptions for comprehensive audio-visual understanding. The term “dense” reflects two key properties: (1) temporally-dense: continuous scene segmentation with explicit timestamps, and (2) description-dense: structured captions spanning six dimensions: Auiod-Visual Events, Visual Background, Camera State, Shot Editing, Dialogue, and Acoustic cues. These “script-like” descriptions allow readers to imagine the video scene-by-scene, as though reading a cinematic screenplay.
@@ 1 Introduction
As video understanding (Xu et al., 2025b; Ye et al., 2025; Yao et al., 2024; Sun et al., 2025, 2024) and generation (Gan et al., 2025) enter the “sound” era, the alignment and interaction of omni-modal information (audio, visual, and text) have become pivotal research directions for Multimodal Large Language Models (MLLMs) (Arefeen et al., 2024; Bai et al., 2025a; Zhang et al., 2024; Yao et al., 2025). Within this context, Omni-Video Captioning, which generates temporally grounded audio-visual-text triplets, emerges as a critical foundational task (Tang et al., 2025; Geng et al., 2025; Yuan et al., 2025a). Such high-quality audio-visual-text data provide comprehensive supervision signals that enable MLLMs to learn fine-grained cross-modal alignment during pre-training and post-training, while also benefiting downstream tasks such as Audio-Visual Reasoning (Zhou et al., 2025; Benchekroun et al., 2023) and Video-to-Audio Generation (Shi et al., 2025).
However, a performant omni-video captioning framework, accompanied by a dedicated benchmark and evaluation suite, remains a largely unexplored frontier in the open-source community. Existing audio-visual captioning (Tang et al., 2025; Wu et al., 2025) works primarily focus on generating global, paragraph-level descriptions without explicit timestamps. This lack of temporal granularity fails to provide the dense supervision signals necessary for MLLMs to master time-aware reasoning, such as temporal grounding (Wang et al., 2025). On the other hand, traditional dense video captioning approaches (Ren et al., 2024; Yang et al., 2023) largely remain confined to the visual modality, neglecting the rich semantics embedded in audio. While recent advanced methods like LongVALE (Geng et al., 2025) have begun to incorporate audio cues, they predominantly focus on identifying salient events and generating concise summaries. This sparse and brief paradigm overlooks the continuous, fine-grained audio-visual nuances, thereby failing to capture the comprehensive semantics required for deep omni-modality alignment.
@@
To facilitate research in this direction, we construct a high-quality benchmark named OmniDCBench, comprising 1,122 human-annotated samples. Evaluating this task presents unique challenges, particularly the ambiguity of continuous scene boundaries. To address this, we propose a novel unified metric SodaM, which jointly measures temporal timestamp accuracy and the semantic completeness of lengthy captions. SodaM incorporates a dynamic programming alignment process to mitigate the time boundary gap between model predictions and human references. Finally, we present a strong baseline TimeChat-Captioner-7B, trained on synthesized high-quality data via Supervised Fine-Tuning (SFT) and Group Relative Policy Optimization (GRPO) stages. Extensive experiments demonstrate that TimeChat-Captioner not only achieves State-of-the-Art performance on OmniDCBench, surpassing Gemini-2.5-Pro (Gemini Team, 2024), but also generates rich semantics that boost performance on downstream Audio-Visual Reasoning tasks like Daily-Omni (Zhou et al., 2025), and WorldSense (Benchekroun et al., 2023), and generalized to temporal grouding task Charades-STA (Gao et al., 2017). We hope TimeChat-Captioner will deliver dense temporal and textual supervision that significantly enhances MLLMs’ omni-modal alignment capabilities.
@@ 2.1 Audio-Visual Captioning
Video captioning aims to generate textual descriptions of video content (Wang et al., 2024; Yuan et al., 2025b), with recent studies exploring fine-grained captioning that describes detailed temporal dynamics (Zhong et al., 2025). The emergence of omni-modal models (Comanici et al., 2025; Xu et al., 2025a; AI et al., 2025) has shifted research from vision-centric to joint audio-visual understanding (Chen et al., 2025). Representative works include AVoCaDO (Chen et al., 2025) for audiovisual temporal coherence, video-SALMONN-2 (Tang et al., 2025), and UGC-VideoCaptioner (Wu et al., 2025) for multimodal integration. However, these methods generate holistic captions without explicit temporal grounding. In contrast, TimeChat-Captioner outputs timestamped captions with structured, fine-grained descriptions for each scene.
@@ 3.1 Task Definition
2) Structured and Fine-grained Output: Whereas exsiting methods typically provide brief descriptions spanning only a few sentences, OmniDenseCaptioning is designed to generate structured, comprehensive narratives across six distinct dimensions. This approach enables the capture of subtle, nuanced visual and audio details, resulting in richer and more informative captions.
@@ 3.3 Evaluation Metric Design
    where $s$ and $e$ denote the start time and end time of each predicted scene, respectively. This handles the common case where MLLMs generate finer-grained (shorter) segments than the ground-truth as Figure 2 (c) shown. We only merge predictions while keeping the ground-truth unchanged to ensure evaluation fairness.
@@ 4.2 Training Data Collection
Video Source Sampling. We curate videos from two complementary datasets: (1) MMTrail-2M (2024), which is donimant and features a diverse and carefully cleaned collection of trailer videos spanning a broad range of topics; and (2) Movie101 (2025), which consists of movie commentary videos with abundant and rich audiovisual content. To balance annotation quality and information density, we segment the raw videos into 3-minute clips.
@@ 4.3.2 GRPO Stage
where each hyperparameters $\alpha$ controls the contribution of its respective reward.
@@ 5.2 Main Results on OmniDCBench
As summarized in Table 1, TimeChat-Captioner achieves highly competitive results on the Omni-Video Scripting Benchmark. Regarding scene boundary localization, our model ranks second only to the industry-leading proprietary model, Gemini-2.5-Pro, while significantly outperforming all other open-source baselines. For time-aware captioning quality, evaluated via the SodaM metric (an aggregate average across six dimensions: camera, events, background, acoustics, shot editing, and dialogue), TimeChat-Captioner-GRPO achieves state-of-the-art (SOTA) performance with a score of 35.0. This result even surpasses the strongest closed-source model, Gemini-2.5-Pro, demonstrating that RL effectively improves accurate scene segmentation and fine-grained captioning. Quantitative cases are visualized in Figure 4.
@@ 6 Conclusions
In future work, we plan to address these limitations through two directions: (1) collecting more diverse long-form videos to improve duration generalization and timestamp segmentation accuracy, and (2) incorporating efficient techniques such as token compression (Yao et al., 2026) to reduce the sequence length of audio-video-text inputs, thereby lowering training costs, especially during the GRPO stage.
@@ Appendix B Additional Experimental Results
Effect of Reward Weights. We investigate the sensitivity of our model to the reward weight coefficients $(\alpha_{f},\alpha_{l},\alpha_{t},\alpha_{c})$ in Equation 8. As shown in Table 5, varying the weight of the coherence reward $R_{C}$ from 1.0 to 1.5 results in marginal performance differences across all metrics (less than 0.5% on F1, mIoU, and SodaM). This suggesting that the four reward components provide complementary supervision signals without requiring extensive hyperparameter tuning.
@@ C.2 Human Annotation Details for OmniDCBench
The OmniDCBench is entirely annotated by human experts. We recruited approximately annotators through crowdsourcing platforms, and the complete annotation process spanned approximately one month. As shown in Figure 7, the annotation interface is designed to be intuitive and user-friendly, featuring clear instructions and real-time feedback mechanisms to facilitate efficient task completion. To ensure annotation quality, each sample was reviewed by at least one additional annotator.

### v2

@@ Abstract
Figure 1: Illustration of the OmniDenseCaptioning task. This paper introduces Omni Dense Captioning task, which generates fine-grained, temporally grounded descriptions for comprehensive audio-visual understanding. The term “dense” reflects two key properties: (1) temporally-dense: continuous scene segmentation with explicit timestamps, and (2) description-dense: structured captions spanning six dimensions: Audio-Visual Events, Visual Background, Camera State, Shot Editing, Dialogue, and Acoustic cues. These “script-like” descriptions allow readers to imagine the video scene-by-scene, as though reading a cinematic screenplay.
@@ 1 Introduction
As video understanding (Xu et al., 2025b; Ye et al., 2025; Yao et al., 2024c; Sun et al., 2025, 2024; Ouyang et al., 2025; Yao et al., 2025b) and generation (Gan et al., 2025) enter the “sound” era, the alignment and interaction of omni-modal information (audio, visual, and text) have become pivotal research directions for Multimodal Large Language Models (MLLMs) (Arefeen et al., 2024; Bai et al., 2025a; Zhang et al., 2024; Yao et al., 2025a). Within this context, Omni-Video Captioning, which generates temporally grounded audio-visual-text triplets, emerges as a critical foundational task (Tang et al., 2025; Geng et al., 2025; Yuan et al., 2025a). Such high-quality audio-visual-text data provide comprehensive supervision signals that enable MLLMs to learn fine-grained cross-modal alignment during pre-training and post-training, while also benefiting downstream tasks such as Audio-Visual Reasoning (Zhou et al., 2025; Benchekroun et al., 2023) and Video-to-Audio Generation (Shi et al., 2025).
However, a performant omni-video captioning framework, accompanied by a dedicated benchmark and evaluation suite, remains a largely unexplored frontier in the open-source community. Existing audio-visual captioning (Tang et al., 2025; Wu et al., 2025) works primarily focus on generating global, paragraph-level descriptions without explicit timestamps. This lack of temporal granularity fails to provide the dense supervision signals necessary for MLLMs to master time-aware reasoning, such as temporal grounding (Wang et al., 2025a). On the other hand, traditional dense video captioning approaches (Ren et al., 2024; Yang et al., 2023) largely remain confined to the visual modality, neglecting the rich semantics embedded in audio. While recent advanced methods like LongVALE (Geng et al., 2025) have begun to incorporate audio cues, they predominantly focus on identifying salient events and generating concise summaries. This sparse and brief paradigm overlooks the continuous, fine-grained audio-visual nuances, thereby failing to capture the comprehensive semantics required for deep omni-modality alignment.
@@
To facilitate research in this direction, we construct a high-quality benchmark named OmniDCBench, comprising 1,122 human-annotated samples. Evaluating this task presents unique challenges, particularly the ambiguity of continuous scene boundaries. To address this, we propose a novel unified metric SodaM, which jointly measures temporal timestamp accuracy and the semantic completeness of lengthy captions. SodaM incorporates a dynamic programming alignment process to mitigate the time boundary gap between model predictions and human references. Finally, we present a strong baseline TimeChat-Captioner-7B, trained on synthesized high-quality data via Supervised Fine-Tuning (SFT) and Group Relative Policy Optimization (GRPO) stages. Extensive experiments demonstrate that TimeChat-Captioner not only achieves State-of-the-Art performance on OmniDCBench, surpassing Gemini-2.5-Pro (Gemini Team, 2024), but also generates rich semantics that boost performance on downstream Audio-Visual Reasoning tasks like Daily-Omni (Zhou et al., 2025), and WorldSense (Benchekroun et al., 2023), and generalized to the temporal grounding task Charades-STA (Gao et al., 2017). We hope TimeChat-Captioner will deliver dense temporal and textual supervision that significantly enhances MLLMs’ omni-modal alignment capabilities.
@@ 2.1 Audio-Visual Captioning
Video captioning aims to generate textual descriptions of video content (Wang et al., 2024a; Yuan et al., 2025b; Wang et al., 2025b, 2024b; Yao et al., 2023, 2024b, 2022), with recent studies exploring fine-grained captioning that describes detailed temporal dynamics (Zhong et al., 2025; Li et al., 2024). The emergence of omni-modal models (Comanici et al., 2025; Xu et al., 2025a; AI et al., 2025) has shifted research from vision-centric to joint audio-visual understanding (Chen et al., 2025). Representative works include AVoCaDO (Chen et al., 2025) for audiovisual temporal coherence, video-SALMONN-2 (Tang et al., 2025), UGC-VideoCaptioner (Wu et al., 2025) for multimodal integration, and DiaDem (Chen et al., 2026) for dialogue-aware audiovisual captioning. However, these methods generate holistic captions without explicit temporal grounding. In contrast, TimeChat-Captioner outputs timestamped captions with structured, fine-grained descriptions for each scene.
@@ 3.1 Task Definition
2) Structured and Fine-grained Output: Whereas existing methods typically provide brief descriptions spanning only a few sentences, OmniDenseCaptioning is designed to generate structured, comprehensive narratives across six distinct dimensions. This approach enables the capture of subtle, nuanced visual and audio details, resulting in richer and more informative captions.
@@ 3.3 Evaluation Metric Design
    where $s$ and $e$ denote the start time and end time of each predicted scene, respectively. This handles the common case where MLLMs generate finer-grained (shorter) segments than the ground-truth as shown in Figure 2 (c). We only merge predictions while keeping the ground-truth unchanged to ensure evaluation fairness.
@@ 4.2 Training Data Collection
Video Source Sampling. We curate videos from two complementary datasets: (1) MMTrail-2M (2024), which is dominant and features a diverse and carefully cleaned collection of trailer videos spanning a broad range of topics; and (2) Movie101 (2025), which consists of movie commentary videos with abundant and rich audiovisual content. To balance annotation quality and information density, we segment the raw videos into 3-minute clips.
@@ 4.3.2 GRPO Stage
where each hyperparameter $\alpha$ controls the contribution of its respective reward.
@@ 5.2 Main Results on OmniDCBench
As summarized in Table 1, TimeChat-Captioner achieves highly competitive results on the Omni-Video Scripting Benchmark. Regarding scene boundary localization, our model ranks second only to the industry-leading proprietary model, Gemini-2.5-Pro, while significantly outperforming all other open-source baselines. For time-aware captioning quality, evaluated via the SodaM metric (an aggregate average across six dimensions: camera, events, background, acoustics, shot editing, and dialogue), TimeChat-Captioner-GRPO achieves state-of-the-art (SOTA) performance with a score of 35.0. This result even surpasses the strongest closed-source model, Gemini-2.5-Pro, demonstrating that RL effectively improves accurate scene segmentation and fine-grained captioning. Qualitative cases are visualized in Figure 4.
@@ 6 Conclusions
In future work, we plan to address these limitations through two directions: (1) collecting more diverse long-form videos to improve duration generalization and timestamp segmentation accuracy, and (2) incorporating efficient techniques such as token compression (Yao et al., 2026, 2024a) to reduce the sequence length of audio-video-text inputs, thereby lowering training costs, especially during the GRPO stage.
@@ Appendix B Additional Experimental Results
Effect of Reward Weights. We investigate the sensitivity of our model to the reward weight coefficients $(\alpha_{f},\alpha_{l},\alpha_{t},\alpha_{c})$ in Equation 8. As shown in Table 5, varying the weight of the coherence reward $R_{C}$ from 1.0 to 1.5 results in marginal performance differences across all metrics (less than 0.5% on F1, mIoU, and SodaM). This suggests that the four reward components provide complementary supervision signals without requiring extensive hyperparameter tuning.
@@ C.2 Human Annotation Details for OmniDCBench
The OmniDCBench is entirely annotated by human experts. We recruited approximately 15 annotators through crowdsourcing platforms, and the complete annotation process spanned approximately one month. As shown in Figure 7, the annotation interface is designed to be intuitive and user-friendly, featuring clear instructions and real-time feedback mechanisms to facilitate efficient task completion. To ensure annotation quality, each sample was reviewed by at least one additional annotator.